OpenStudy (howard-wolowitz):

14. What is the range of the following data set? 7.7, 8.4, 9, 8, 6.9 0.4 0.8 1.4 2.1

OpenStudy (rational):

range = maximum - minimum

OpenStudy (howard-wolowitz):

15. Assume that the weight loss for the first month of a diet program varies between 6 pounds and 12 pounds, and is spread evenly over the range of possibilities, so that there is a uniform distribution. Find the probability that the weight lost is less than 11 pounds A. 1/6 B. 5/6 C. 5/7 D. 1/3

OpenStudy (howard-wolowitz):

but for 14 wouldnt it be 2.1?

OpenStudy (rational):

Yes!

OpenStudy (howard-wolowitz):

oh i see now ok gotcha

OpenStudy (howard-wolowitz):

16. Given a mean of 8 and a standard deviation of 0.7, what is the z-score of the value 9 rounded to the nearest tenth? A. -1.4 B. -1.0 C. 1.0 D. 1.4

OpenStudy (howard-wolowitz):

17. Provide an appropriate response. Find the area of the shaded region. The graph depicts IQ scores of adults, and those scores are normally distributed with a mean of 100 and a standard deviation of 15 (as on the Wechsler test).

OpenStudy (howard-wolowitz):

17. 0.7486 0.6293 0.4400 0.8051

OpenStudy (howard-wolowitz):

18. Provide an appropriate response. Find the indicated IQ score. The graph depicts IQ scores of adults, and those scores are normally distributed with a mean of 100 and a standard deviation of 15 (as on the Wechsler test) 119.2 108.1 100.5 80.8

OpenStudy (howard-wolowitz):

18.

OpenStudy (howard-wolowitz):

19. Given a mean of 9 and a standard deviation of 2.8, what is the z-score of the value 11 rounded to the nearest tenth?

OpenStudy (howard-wolowitz):

19. A. -0.7 B. 0.7 C. 0.9 D. -0.9

OpenStudy (rational):

use this formula : \[\text{z score} = \dfrac{\text{observation - mean} }{\text{standard deviation}}\]

OpenStudy (rational):

\[\text{z score} = \dfrac{\text{11 - 9} }{\text{2.8}} = ?\]

OpenStudy (howard-wolowitz):

solve for z?

OpenStudy (howard-wolowitz):

i got a deciaml

OpenStudy (rational):

\[\dfrac{\text{11 - 9} }{\text{2.8}} = ?\]

OpenStudy (rational):

what did u get ?

OpenStudy (howard-wolowitz):

0.71428

OpenStudy (rational):

thats right! rounding to nearest tenth : \[0.71428\approx 0.7\]

OpenStudy (howard-wolowitz):

20. Find the indicated z score. The graph depicts the standard normal distribution with mean 0 and standard deviation 1. Shaded area is 0.4013

OpenStudy (howard-wolowitz):

20.

OpenStudy (howard-wolowitz):

20. A. -0.25 B. -0.57 C. 0.57 D. 0.25

OpenStudy (howard-wolowitz):

3. If z is a standard normal variable, find the probability. The probability that z is less than 1.13 0.8485 0.8907 0.1292 0.8708

OpenStudy (howard-wolowitz):

if you dont mind i have about 4 already answered i just wont you to check

OpenStudy (howard-wolowitz):

3 being one of them

OpenStudy (rational):

0.8708 http://www.wolframalpha.com/input/?i=P%28X%3C1.13%29

OpenStudy (howard-wolowitz):

4. A set of data with a mean of 45 and a standard deviation of 8.3 is normally distributed. Find the value that is –2 standard deviations away from the mean. 20.1 43.0 61.6 28.4

OpenStudy (howard-wolowitz):

-2 where it blurs it

OpenStudy (howard-wolowitz):

see one guy did these earlier and got something else then what u r getting

OpenStudy (howard-wolowitz):

5. A set of data with a mean of 45 and a standard deviation of 8.3 is normally distributed. Find the value that is +1 standard deviation away from the mean. 36.7 28.4 53.3 45.0

OpenStudy (howard-wolowitz):

6. A set of data with a mean of 45 and a standard deviation of 8.3 is normally distributed. Find the value that is +3 standard deviations away from the mean. 48.0 20.1 78.9 69.9

OpenStudy (rational):

we already did this in previous post right ?

OpenStudy (howard-wolowitz):

oh ok

OpenStudy (howard-wolowitz):

i forgot

OpenStudy (howard-wolowitz):

2. What is the standard deviation of the following data set rounded to the nearest tenth? 300, 785, 670, 187, 760, 724 571 570 259.1 236.6

OpenStudy (howard-wolowitz):

for this one i think its c or d

OpenStudy (rational):

d is correct

OpenStudy (howard-wolowitz):

cool well you were a lot of help and if you wont for more medals i have another 15 i have to do if u wanna help... up to u though...... THANKS

OpenStudy (rational):

post as a separate thread and tag me, il try..